Sweeten Company had no jobs in progress at the beginning of March and no beginning inventories. It started only two jobs during March—Job P and Job Q. Job P was completed and sold by the end of March and Job Q was incomplete at the end of March. The company uses a plantwide predetermined overhead rate based on direct labor-hours. The following additional information is available for the company as a whole and for Jobs P and Q (all data and questions relate to the month of March):

Estimated total fixed manufacturing overhead $ 13,200
Estimated variable manufacturing overhead per direct labor-hour $ 1.20
Estimated total direct labor-hours to be worked 3,300
Total actual manufacturing overhead costs incurred $ 17,000


Job P Job Q
Direct materials $ 13,100 $ 9,300
Direct labor cost $ 43,200 $ 11,700
Actual direct labor-hours worked 2,400 650

6. What is the amount of underapplied or overapplied overhead?
Answer #1

Predetermine overhead rate = 13200/3300+1.20 = 5.20 per DLH

Applied overhead = (2400+650)*5.20 = 15860

Actual overhead = 17000

Under applied overhead = 15860-17000 = 1140